OpenStudy (anonymous):

Help Please, give a medal and fan to first that answers In the below system, solve for y in the first equation. 3x + y = 9 4x - 3y = 6

OpenStudy (anonymous):

@JamesJ

OpenStudy (anonymous):

already got the answer it is -3x + 9

OpenStudy (raden_zaikaria):

3 (3x + y ) =3( 9) is 9x + 3y = 27

OpenStudy (anonymous):

i already got the answer

OpenStudy (raden_zaikaria):

ok hehe i want to try

OpenStudy (anonymous):

wow

OpenStudy (raden_zaikaria):

( 4x - 3y ) + (9x + 3y) = 6 - 27

OpenStudy (raden_zaikaria):

( 4x - 3y ) - (9x + 3y) = 6 + 27 this is correct haiya hehe

OpenStudy (anonymous):

the answer is -3x+9

OpenStudy (raden_zaikaria):

-5x = 33

OpenStudy (anonymous):

a. -3x + 9 b. 3x - 9 c. 3x + 9 d. -3x - 9 these are the choices

OpenStudy (raden_zaikaria):

0ic hehehe ur give a choice hehehe